A ball is dropped from rest from a height of 10.0 m. Calculate:
a) The time it takes to hit the ground and the speed with which it reaches it.
b) The speed it has right after the first boat, if with the rebound it loses 20% of the initial speed.
c) The maximum height of the second boat.
1
Expert's answer
2020-02-12T10:48:30-0500

a)


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(10)}{9.8}}=1.43\ s"

"v=\sqrt{2gh}=\sqrt{2(9.8)(10)}=14\frac{m}{s}"

b)


"v'=v(1-0.2)=14(0.8)=11.2\frac{m}{s}"

c)


"h'=\frac{v'^2}{2g}=\frac{11.2^2}{2(9.8)}=6.4\ m"